# Vizing’s Theorem in Near-Linear Time

### v1

@@ 32             $S\leftarrow S\setminus P_{\operatorname*{\mathbf{F}}^{\prime}}$
[…]
@@
The subroutine Reduce-U-Edges satisfies 5.5.
@@
[…]
@@
[…]
@@
(1) If we enter the if statement on Algorithm 4, then calling $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$ and $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}}^{\prime})$ after uncoloring edges on their Vizing chains does not damage any u-component in $\operatorname*{\mathcal{U}}$ apart from $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}$ and $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}^{\prime}}$, which are both removed from $\operatorname*{\mathcal{U}}$. This is because shifting colors around the Vizing fans $\operatorname*{\mathbf{F}}$ and $\operatorname*{\mathbf{F}}^{\prime}$ does not damage any u-components apart from $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}$ and $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}^{\prime}}$ (see Lemma 3.9) and truncating an alternating path before flipping it ensures that we do not remove any colors from the palettes of the vertices on that path. Finally, note that the u-fan $\operatorname*{\boldsymbol{f}}$ that we add to $\operatorname*{\mathcal{U}}$ only uses colors that were previously unavailable. (2) If we enter the else statement on Algorithm 4, then the argument is completely analogous to the previous case, except that we do not add any u-components to $\operatorname*{\mathcal{U}}$ (see Lemma 3.9). (3) If we enter the if statement on Algorithm 4, then calling $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$ damages at most one u-component in $\operatorname*{\mathcal{U}}$ apart from $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}$ (see Lemma 3.9). In particular, if there is such a u-component $\operatorname*{\boldsymbol{g}}$, it will contain the vertex $y$ and have $c_{\operatorname*{\boldsymbol{g}}}(x)=\chi(x,y)$. Thus, we check if such a u-component exists and remove it from $\operatorname*{\mathcal{U}}$.
@@
[…]
@@
The subroutine Reduce-U-Edges satisfies 5.6.
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Let $\mathcal{F}^{\star}$ denote the set $\mathcal{F}$ when we first call the subroutine Reduce-U-Edges. For each $\operatorname*{\mathbf{F}}\in\mathcal{F}^{\star}$, let $L_{\operatorname*{\mathbf{F}}}$ denote the value of $L$ at the time that $\operatorname*{\mathbf{F}}$ is removed from $\mathcal{F}$.
@@
[…]
@@
Now, assume w.l.o.g. that =$\mathcal{F}^{\star}=\{\operatorname*{\mathbf{F}}_{1},\dots,\operatorname*{\mathbf{F}}_{\lambda_{\alpha}}\}$ and $L_{\lambda_{\alpha}}\leq\dots\leq L_{1}$, where $L_{i}:=L_{\operatorname*{\mathbf{F}}_{i}}$. Recall that $m_{\alpha}$ denotes the number of edges with color $\alpha$ when we first call the subroutine.
@@
For each $i\in[\lambda_{\alpha}]$, we have that $L_{i}\leq 4(m_{\alpha}+2\lambda_{\alpha})/i$.
@@
Let $i\in[\lambda_{\alpha}]$, $\tau=4(m_{\alpha}+2\lambda_{\alpha})/i$ and suppose that $\tau<L_{i}$. Then, at the start of iteration $\tau+1$, the Vizing fans $\operatorname*{\mathbf{F}}_{1},\dots\operatorname*{\mathbf{F}}_{i}$ are all contained in $\mathcal{F}$. Furthermore, it follows from Invariants 5.5 and 5.6 that, at the start of iteration $\tau+1$, the alternating paths $P_{\operatorname*{\mathbf{F}}_{1}},\dots,P_{\operatorname*{\mathbf{F}}_{i}}$ form a collection of $i$ edge-disjoint $\{\alpha,\cdot\}$-alternating paths of length $\tau$ in $G$. Let $T$ denote the total length of these paths. We can observe that $T$ is at most $3m_{\alpha}^{\prime}$, where $m_{\alpha}^{\prime}$ is the number of edges that currently have color $\alpha$, since at least a third of the edges in each of these paths has color $\alpha$ (note that any $\{\alpha,\cdot\}$-alternating path of length $k$ has at least $\lfloor k/2\rfloor$ edges with color $\alpha$). We can also observe that $m_{\alpha}^{\prime}\leq m_{\alpha}+2\lambda_{\alpha}$ since the number of edges with color $\alpha$ increases by at most $2$ each time we activate a Vizing fan in $\mathcal{F}$. Thus, it follows that
$4(m_{\alpha}+2\lambda_{\alpha})=i\tau\leq T\leq 3m_{\alpha}^{\prime}\leq 3(m_{\alpha}+2\lambda_{\alpha}),$
giving a contradiction. ∎
@@
We can implement Reduce-U-Edges to run in $O((m_{\alpha}+\Delta\lambda_{\alpha})\log n)$ time.
@@
Activating the trivial Vizing fans in $\mathcal{F}$ when the subroutine is first called can be done in $O(\Delta\lambda_{\alpha})$ time. The running time of the rest of the subroutine is dominated by the time taken to handle the call to Update-Path. Combining Claims 5.11 and 5.12, it follows that the total time spent handling calls to Update-Path is at most
$O(\Delta\lambda_{\alpha})+O(1)\cdot\sum_{\operatorname*{\mathbf{F}}\in\mathcal{F}^{\star}}L_{\operatorname*{\mathbf{F}}}\leq O(\Delta\lambda_{\alpha})+O(1)\cdot\sum_{i=1}^{\lambda_{\alpha}}\frac{m_{\alpha}+\lambda_{\alpha}}{i}$
$\leq O((m_{\alpha}+\Delta\lambda_{\alpha})\log n).\qed$
5.3 Analysis of Construct-U-Fans (Proof of Lemma 4.1)
@@
The algorithm first initializes an empty separable collection $\operatorname*{\mathcal{U}}=\varnothing$. It then scans through the edges of the graph and retrieves the $\lambda$ edges $e_{1},\dots,e_{\lambda}$ that are uncolored by $\chi$. The algorithm then scans through each of these uncolored edges and, for each $e_{i}=(u_{i},v_{i})$, picks a missing color $\alpha_{i}\in\mathsf{miss}_{\chi}(u_{i})\setminus C_{\operatorname*{\mathcal{U}}}(u_{i})$ (see 3.5) and adds the u-edge $\operatorname*{\boldsymbol{e}}_{i}:=(u_{i},v_{i},\alpha_{i})$ to $\operatorname*{\mathcal{U}}$. It’s easy to verify that the resulting collection $\operatorname*{\mathcal{U}}$ is separable and contains $\lambda$ u-edges. Using the data structures outlined in Section 3.2, finding such a color and adding a u-edge to $\operatorname*{\mathcal{U}}$ can be done in $O(1)$ time. Thus, this entire process can be implemented in $O(m)$ time. ∎
@@
Let $\lambda_{\alpha}$ denote the number of $\alpha$-primed u-edges in the initial collection $\operatorname*{\mathcal{U}}$. Note that the number of such u-edges can only decrease throughout the run of this algorithm. It follows from Lemmas 5.4 and 5.13 that the total running time of the algorithm Construct-U-Fans across all of these calls to the subroutines Prune-Vizing-Fans and Reduce-U-Edges is at most
$\sum_{\alpha\in[\Delta+1]}\left(O(\lambda_{\alpha}\Delta)+O((m_{\alpha}+\lambda+\lambda_{\alpha}\Delta)\log n)\right)\leq O((m+\Delta\lambda)\log n).$
@@
The algorithm Construct-U-Fans either extends the coloring $\chi$ to at least $\lambda/9$ more edges or returns a separable collection of at least $\lambda/9$ u-fans $\operatorname*{\mathcal{U}}$.
@@
To see why this lemma holds, consider the following three quantities and how they evolve over time throughout the execution of the algorithm Construct-U-Fans: The number of edges that the algorithm has extended the coloring to so far, $\Psi_{\texttt{c}}$, the number of u-edges in $\operatorname*{\mathcal{U}}$, $\Psi_{\texttt{e}}:=|\mathcal{E}|$, and the number of u-fans in $\operatorname*{\mathcal{U}}$, $\Psi_{\texttt{f}}:=|\mathcal{U}\setminus\mathcal{E}|$. Immediately after constructing the separable collection of $\lambda$ u-edges $\operatorname*{\mathcal{U}}$, it holds that $\Psi_{\texttt{c}}=0$, $\Psi_{\texttt{e}}=\lambda$ and $\Psi_{\texttt{f}}=0$. We next argue that by the time the execution of the algorithm has finished, either $\Psi_{\texttt{c}}\geq\lambda/9$ or $\Psi_{\texttt{f}}\geq\lambda/9$ must hold. To this end, we employ the following potential function argument.
@@
Immediately after constructing the separable collection of $\lambda$ u-edges in $\operatorname*{\mathcal{U}}$, we have $3(\Psi_{\texttt{f}}+2\Psi_{\texttt{c}})+\Psi_{\texttt{e}}=\lambda$. 5.17 implies that $3(\Psi_{\texttt{f}}+2\Psi_{\texttt{c}})+\Psi_{\texttt{e}}\geq\lambda$ holds at all times afterwards. Since the algorithm iterates until $\mathcal{E}=\varnothing$, we know that $\Psi_{\texttt{e}}=0$ when the algorithm’s execution terminates, so at that moment we have $3(\Psi_{\texttt{f}}+2\Psi_{\texttt{c}})\geq\lambda$. Hence, either $\Psi_{\texttt{f}}\geq\lambda/9$ or $\Psi_{\texttt{c}}\geq\lambda/9$ must hold. ∎
6 The Algorithm Color-U-Fans (Proof of Lemma 4.2)

### v3

@@ 32    $S\leftarrow S\setminus P_{\operatorname*{\mathbf{F}}^{\prime}}$
Algorithm 5 $\textnormal{{Update-Path}}(\operatorname*{\mathbf{F}})$
@@
The subroutine Reduce-U-Edges satisfies 7.5.
@@
We show that, as long as 7.5 is satisfied, calling $\textnormal{{Update-Path}}(\operatorname*{\mathbf{F}})$ for some $\operatorname*{\mathbf{F}}\in\mathcal{F}$ maintains the invariant. We first note that 7.5 is trivially satisfied immediately after initializing $L$, $S$ and $\{P_{\operatorname*{\mathbf{F}}}\}_{\operatorname*{\mathbf{F}}\in\mathcal{F}}$.
@@
Now, assume that 7.5 is satisfied and suppose that we call $\textnormal{{Update-Path}}(\operatorname*{\mathbf{F}})$ for some $\operatorname*{\mathbf{F}}\in\mathcal{F}$. We now argue that each condition of 7.5 is satisfied after handling this call.
@@
  1. (1)
    If we enter the if statement on Algorithm 5, then calling $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$ and $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}}^{\prime})$ after uncoloring edges on their Vizing chains does not damage any u-component in $\operatorname*{\mathcal{U}}$ apart from $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}$ and $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}^{\prime}}$, which are both removed from $\operatorname*{\mathcal{U}}$. This is because shifting colors around the Vizing fans $\operatorname*{\mathbf{F}}$ and $\operatorname*{\mathbf{F}}^{\prime}$ does not damage any u-components apart from $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}$ and $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}^{\prime}}$ (see Lemma 5.9) and truncating an alternating path before flipping it ensures that we do not remove any colors from the palettes of the vertices on that path. Finally, note that the u-fan $\operatorname*{\boldsymbol{f}}$ that we add to $\operatorname*{\mathcal{U}}$ only uses colors that were previously unavailable. See the example Figure 6 for an illustration.
[…]
    Figure 6: In this picture, $\alpha$ is blue, and two uncolored edges $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}=(u_{1},v_{1}),\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}^{\prime}}=(u_{2},v_{2})$ generate Vizing fans $\operatorname*{\mathbf{F}}$ and $\operatorname*{\mathbf{F}}^{\prime}$, and the two Vizing chains are joining at $(x,y)$ for the first time and in the same direction. Then, we can rotate both Vizing fans and partially flip $P_{\operatorname*{\mathbf{F}}}$ and $P_{\operatorname*{\mathbf{F}}^{\prime}}$ to make a u-fan
  2. (2)
    If we enter the else statement on Algorithm 5, then the argument is completely analogous to the previous case, except that we do not add any u-components to $\operatorname*{\mathcal{U}}$ (see Lemma 5.9). See Figure 7 for an illustration.
[…]
    Figure 7: In this picture, $\alpha$ is blue, and two uncolored edges $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}=(u_{1},v_{1}),\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}^{\prime}}=(u_{2},v_{2})$ generate Vizing fans $\operatorname*{\mathbf{F}}$ and $\operatorname*{\mathbf{F}}^{\prime}$, and the two Vizing chains are joining at $(x,y)$ for the first time and in the opposite direction. Then, we can uncolor $(x,y)$ and color both $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}},\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}^{\prime}}$.
  3. (3)
    If we enter the if statement on Algorithm 5, then calling $\textnormal{{Vizing}}(\operatorname*{\mathbf{F}})$ damages at most one u-component in $\operatorname*{\mathcal{U}}$ apart from $\operatorname*{\boldsymbol{e}}_{\operatorname*{\mathbf{F}}}$ (see Lemma 5.9). In particular, if there is such a u-component $\operatorname*{\boldsymbol{g}}$, it will contain the vertex $y$ and have $c_{\operatorname*{\boldsymbol{g}}}(x)=\chi(x,y)$. Thus, we check if such a u-component exists and remove it from $\operatorname*{\mathcal{U}}$.
@@
Condition 1: Let $\operatorname*{\mathbf{F}}^{\prime}$ be some Vizing fan that remains in $\mathcal{F}$ after handling this call. We now show that the first $|P_{\operatorname*{\mathbf{F}}^{\prime}}|$ edges of $\textnormal{{Vizing-Path}}(\operatorname*{\mathbf{F}}^{\prime})$ do not change throughout this call. It follows from the argument for Condition 5 that neither the colors of the edges nor the palettes of the vertices in $\operatorname*{\mathbf{F}}^{\prime}$ change during this call. It remains to show that none of the edges in $P_{\operatorname*{\mathbf{F}}^{\prime}}$ change color during this call. We can see that $P_{\operatorname*{\mathbf{F}}^{\prime}}$ and $P_{\operatorname*{\mathbf{F}}}$ are edge-disjoint since otherwise $\operatorname*{\mathbf{F}}^{\prime}$ would have been removed from $\mathcal{F}$. Consequently, $P_{\operatorname*{\mathbf{F}}^{\prime}}$ is edge-disjoint from any Vizing chain that is activated during the call, and hence none of the edges in $P_{\operatorname*{\mathbf{F}}^{\prime}}$ change colors. ∎
@@
The subroutine Reduce-U-Edges satisfies 7.6.
@@
This follows from the fact that initially $|P_{\operatorname*{\mathbf{F}}}|=0$ for all $\operatorname*{\mathbf{F}}\in\mathcal{F}$ and that the subroutine calls $\textnormal{{Update-Path}}(\operatorname*{\mathbf{F}})$ for each $\operatorname*{\mathbf{F}}\in\mathcal{F}$ during each round, which increases $|P_{\operatorname*{\mathbf{F}}}|$ by $1$. ∎
@@
While activating some trivial Vizing fan $\operatorname*{\mathbf{F}}_{\operatorname*{\boldsymbol{e}}}\in\mathcal{F}$ at the start, we remove $\operatorname*{\boldsymbol{e}}$ from $\operatorname*{\mathcal{U}}$ (and thus also from $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$) along with at most one other u-component in $\operatorname*{\mathcal{U}}$ that is damaged by this operation (see Lemma 5.9). Now, suppose that the subroutine calls $\textnormal{{Update-Path}}(\operatorname*{\mathbf{F}})$ for some $\operatorname*{\mathbf{F}}\in\mathcal{F}$. By considering each case, we can verify that the subroutine either makes no changes to $\chi$ and $\operatorname*{\mathcal{U}}$ or it removes at most $2$ u-edges from $\operatorname*{\mathcal{E}}_{\alpha}(\operatorname*{\mathcal{U}})$ and either (1) extends the coloring $\chi$ to one more edge (see Lines 5-5 and Algorithm 5) and removes at most one other u-component from $\operatorname*{\mathcal{U}}$ (see Algorithm 5), or (2) adds a u-fan to $\operatorname*{\mathcal{U}}$ (see Algorithm 5). ∎
@@
Each call to $\textnormal{{Update-Path}}(\operatorname*{\mathbf{F}})$ for some $\operatorname*{\mathbf{F}}\in\mathcal{F}$ that does not remove $\operatorname*{\mathbf{F}}$ from $\mathcal{F}$ takes $O(1)$ time. Otherwise, it takes $O(\Delta+L)$ time.
@@
By Invariants 7.5 and 7.6, we know that $P_{\operatorname*{\mathbf{F}}}=\textnormal{{Vizing-Path}}(\operatorname*{\mathbf{F}})_{\leq L}$ when the subroutine calls $\textnormal{{Update-Path}}(\operatorname*{\mathbf{F}})$. Thus, updating $P_{\operatorname*{\mathbf{F}}}$ to $\textnormal{{Vizing-Path}}(\operatorname*{\mathbf{F}})_{\leq L+1}$ only requires computing the next edge in the path, which can be done in $O(1)$ time using our data structures (see Section 9).
@@
Let $\mathcal{F}^{\star}$ denote the subset of Vizing fans that get removed from $\mathcal{F}$ by the subroutine Reduce-U-Edges before it terminates. For each $\operatorname*{\mathbf{F}}\in\mathcal{F}^{\star}$, let $L_{\operatorname*{\mathbf{F}}}$ denote the value of $L$ at the time that $\operatorname*{\mathbf{F}}$ is removed from $\mathcal{F}$.
@@
Let $\operatorname*{\mathbf{F}}\in\mathcal{F}^{\star}$. We can observe that the subroutine calls $\textnormal{{Update-Path}}(\operatorname*{\mathbf{F}})$ at most $L_{\operatorname*{\mathbf{F}}}$ times. It follows from 7.10 that the last call takes $O(\Delta+L_{\operatorname*{\mathbf{F}}})$ time while the rest take $O(1)$ time. Thus, the total time spent handling calls to $\textnormal{{Update-Path}}(\operatorname*{\mathbf{F}})$ is at most $O(\Delta+L_{\operatorname*{\mathbf{F}}})$. Summing over each $\operatorname*{\mathbf{F}}\in\mathcal{F}^{\star}$, we get that the total time spent handling calls to Update-Path is at most
@@
Recall that $m_{\alpha}$ denotes the number of edges with color $\alpha$ when we first call the subroutine.
@@
For each $\operatorname*{\mathbf{F}}\in\mathcal{F}^{\star}$, we have that $L_{\operatorname*{\mathbf{F}}}\leq O(m_{\alpha}/\lambda_{\alpha})$.
@@
Let $L_{\max}$ denote the value of $L$ at the start of the final round performed by the subroutine. We now show that $L_{\max}\leq 8(m_{\alpha}+4)/\lambda_{\alpha}$, which implies the claim. At the start of the final round, we know that $|\mathcal{F}|>\lambda_{\alpha}/2$, otherwise the subroutine would terminate. Furthermore, it follows from Invariants 7.5 and 7.6 that, at the start of this round, the alternating paths $\{P_{\operatorname*{\mathbf{F}}}\}_{\operatorname*{\mathbf{F}}\in\mathcal{F}}$ form a collection of at least $\lambda_{\alpha}/2$ edge-disjoint $\{\alpha,\cdot\}$-alternating paths of length $L_{\max}$ in $G$. Let $T$ denote the total length of these paths. We can observe that $T$ is at most $3m_{\alpha}^{\prime}$, where $m_{\alpha}^{\prime}$ is the number of edges that currently have color $\alpha$, since at least a third of the edges in each of these paths has color $\alpha$ (note that any $\{\alpha,\cdot\}$-alternating path of length $k$ has at least $\lfloor k/2\rfloor$ edges with color $\alpha$). We can also observe that $m_{\alpha}^{\prime}\leq m_{\alpha}+2\lambda_{\alpha}$ since the number of edges with color $\alpha$ increases by at most $2$ each time we activate a Vizing fan in $\mathcal{F}$. Thus, it follows that
$\frac{\lambda_{\alpha}}{2}\cdot L_{\max}\leq T\leq 3m_{\alpha}^{\prime}\leq 3(m_{\alpha}+2\lambda_{\alpha}),$
and so $L_{\max}\leq 6m_{\alpha}/\lambda_{\alpha}+12$. ∎
@@
We can implement Reduce-U-Edges to run in $O(m_{\alpha}+\Delta\lambda_{\alpha})$ time.
@@
Activating the trivial Vizing fans in $\mathcal{F}$ when the subroutine is first called can be done in $O(\Delta\lambda_{\alpha})$ time. The running time of the rest of the subroutine is dominated by the time taken to handle the call to Update-Path. Combining Claims 7.11 and 7.12, it follows that the total time spent handling calls to Update-Path is at most
$O(\Delta\lambda_{\alpha})+O(1)\cdot\sum_{\operatorname*{\mathbf{F}}\in\mathcal{F}^{\star}}L_{\operatorname*{\mathbf{F}}}\leq O(\Delta\lambda_{\alpha})+O(\lambda_{\alpha})\cdot O\!\left(\frac{m_{\alpha}}{\lambda_{\alpha}}\right)\leq O(m_{\alpha}+\Delta\lambda_{\alpha}).\qed$
7.3 Analysis of Construct-U-Fans: Proof of Lemma 6.2
@@
The algorithm first initializes an empty separable collection $\operatorname*{\mathcal{U}}=\varnothing$. It then scans through the edges of the graph and retrieves the $\lambda$ edges $e_{1},\dots,e_{\lambda}$ that are uncolored by $\chi$. The algorithm then scans through each of these uncolored edges and, for each $e_{i}=(u_{i},v_{i})$, picks a missing color $\alpha_{i}\in\mathsf{miss}_{\chi}(u_{i})\setminus C_{\operatorname*{\mathcal{U}}}(u_{i})$ (see 5.5) and adds the u-edge $\operatorname*{\boldsymbol{e}}_{i}:=(u_{i},v_{i},\alpha_{i})$ to $\operatorname*{\mathcal{U}}$. It’s easy to verify that the resulting collection $\operatorname*{\mathcal{U}}$ is separable and contains $\lambda$ u-edges. Using the data structures outlined in Section 5.2, finding such a color and adding a u-edge to $\operatorname*{\mathcal{U}}$ can be done in $O(1)$ time. Thus, this entire process can be implemented in $O(m)$ time. ∎
@@
Let $\lambda_{\alpha}$ denote the number of $\alpha$-primed u-edges in the initial collection $\operatorname*{\mathcal{U}}$. Note that the number of such u-edges can only decrease throughout the run of this algorithm. It follows from Lemmas 7.4 and 7.13 that the total running time of the algorithm Construct-U-Fans across all of these calls to the subroutines Prune-Vizing-Fans and Reduce-U-Edges is at most
$\sum_{\alpha\in[\Delta+1]}\left(O(\lambda_{\alpha}\Delta)+O(m_{\alpha}+\lambda+\lambda_{\alpha}\Delta)\right)\leq O(m+\Delta\lambda).$
@@
The algorithm Construct-U-Fans either extends the coloring $\chi$ to at least $\lambda/18$ more edges or returns a separable collection of at least $\lambda/18$ u-fans $\operatorname*{\mathcal{U}}$.
@@
To see why this lemma holds, consider the following three quantities and how they evolve over time throughout the execution of the algorithm Construct-U-Fans: The number of edges that the algorithm has extended the coloring to so far, $\Psi_{\texttt{c}}$, the number of u-edges in $\operatorname*{\mathcal{U}}$, $\Psi_{\texttt{e}}:=|\mathcal{E}|$, and the number of u-fans in $\operatorname*{\mathcal{U}}$, $\Psi_{\texttt{f}}:=|\mathcal{U}\setminus\mathcal{E}|$. Immediately after constructing the separable collection of $\lambda$ u-edges $\operatorname*{\mathcal{U}}$, it holds that $\Psi_{\texttt{c}}=0$, $\Psi_{\texttt{e}}=\lambda$ and $\Psi_{\texttt{f}}=0$. We next argue that by the time the execution of the algorithm has finished, either $\Psi_{\texttt{c}}\geq\lambda/18$ or $\Psi_{\texttt{f}}\geq\lambda/18$ must hold. To this end, we employ the following potential function argument.
@@
Immediately after constructing the separable collection of $\lambda$ u-edges in $\operatorname*{\mathcal{U}}$, we have $3(\Psi_{\texttt{f}}+2\Psi_{\texttt{c}})+\Psi_{\texttt{e}}=\lambda$. 7.17 implies that $3(\Psi_{\texttt{f}}+2\Psi_{\texttt{c}})+\Psi_{\texttt{e}}\geq\lambda$ holds at all times afterwards. Since we have that $|\mathcal{E}|\leq\lambda/2$ when the algorithm terminates, we know that $\Psi_{\texttt{e}}\leq\lambda/2$ at this time, so at that moment we have $3(\Psi_{\texttt{f}}+2\Psi_{\texttt{c}})\geq\lambda/2$. Hence, either $\Psi_{\texttt{f}}\geq\lambda/18$ or $\Psi_{\texttt{c}}\geq\lambda/18$ must hold. ∎
8 The Algorithm Color-U-Fans: Proof of Lemma 6.3